# Magnetic properties of monolayer, multilayer, and bulk CrTe2

### v2

@@ Abstract
We investigate magnetic properties of CrTe2 within the density functional theory (DFT) approach in ferromagnetic phase and combination of DFT and dynamical mean field theory (DFT+DMFT) approach in paramagnetic phase. We show that few layer CrTe2 possesses well formed local magnetic moments. In the single layer CrTe2 we find antiferromagnetic exchange with 120°$\degree$ antiferromagnetic structure most preferable. In the bilayer and trilayer systems electronic correlations in DFT+DMFT approach yield ferromagnetic exchange interaction within each layer, but the interaction between the layers is antiferromagnetic, such that alternation of the direction of magnetization of the layers is expected. In bulk CrTe2 we find the tendency to ferromagnetic order at low temperature, but with increase of temperature antiferromagnetic correlations between the layers dominate. Determination of the critical number of layers at which the interlayer antiferromagnetic order changes to the ferromagnetic one, requires consideration of the non-local Coulomb interactions.
@@ I Introduction
Atomically thin two-dimensional (2D) van der Waals (vdW) materials have attracted significant attention over the past decade due to their exceptional properties Rev1 ; Rev2 ; Rev3 ; Rev4 ; Rev5 ; Rev6 . In particular, they can posses large magnetic moments, substantial Curie temperature, comparable to room temperature, and strong anisotropy. With these properties, 2D magnetic materials are promising for using in the fields of spintronics Rev2 and memory storage devices Rev5 . Currently, many 2D ferromagnetic (FM) materials have been reported; however, a key challenge remains: achieving ferromagnetism at room temperature. According to the Mermin–Wagner theorem, long-range magnetic order cannot exist in two dimensions in isotropic systems at finite temperatures due to thermal fluctuations. However, it is possible to circumvent this limitation by magnetic anisotropy.
@@
Recently, chromium ditelluride CrTe2 has been suggested to exhibit ferromagnetic order even up to room temperature, making it an attractive candidate for further research in this field. Bulk CrTe2 was shown freitas_ferromagnetism_2015 to posses FM order with $T_{C}=330$ K and in-plane magnetic anisotropy. Recent years several groups obtained thin CrTe2 flakes via mechanical exfoliation purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , chemical vapor deposition meng_anomalous_2021 and molecular beam epitaxy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; liu_wafer-scale_2023 ; zheng_two-dimensional_2023 , showing stable FM order with $T_{C}\sim 200$ K. While thick films were shown to have an in-plane magnetic anisotropy purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , the films thinner 10 monolayers show the out-of-plane magnetic anisotropy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; meng_anomalous_2021 . The authors of Refs. zhang_room-temperature_2021 ; meng_anomalous_2021 ; xian_spin_2022 ; liu_wafer-scale_2023 ; wang_strain-_2024 succeeded fabricating monolayer devices. At the same time, the results of the magnetic state of monolayers remain controversial. Experimentally both, FM zhang_room-temperature_2021 ; meng_anomalous_2021 ; liu_wafer-scale_2023 and zigzag antiferromagnetic (AFM-zz) xian_spin_2022 states were reported.
@@
Theoretically, the density functional theory (DFT) approach was used to study various magnetic orders and exchange interactions in CrTe2. In particular, FM lv_strain-controlled_2015 ; li_tunable_2021 ; liu_structural_2022 ; yang_tunable_2021 ; yao_control_2023 , AFM gao_thickness_2021 ; AFM2 ; abuawwad_noncollinear_2022 , charge density wave otero_fumega_controlled_2020 and incommensurate abuawwad_noncollinear_2022 states were considered as possible candidates in single layer CrTe2. It was suggested in Ref. gao_thickness_2021 that the the intralayer coupling of 2 to 4 layer CrTe2 is still antiferromagnetic, similar to that of monolayer CrTe2, and changes to ferromagnetic for larger number of layers. At the same time, the interlayer exchange interaction was also found to change from AFM to FM between 4 and 5 layers. The change of the magnetic orders was related by authors of Ref. gao_thickness_2021 to the change of lattice constants from single-layer to the bulk. It was suggested in Ref. wu_-plane_2022 ; zhu_insight_2023 ; gao_thickness_2021 ; lv_strain-controlled_2015 that relatively small strain can cause transition from AFM-zz to FM what explains experimental data and theoretical discrepancies.
@@
At the same time, Cr is an open almost half filled $d$ shell metal and possesses strong correlations. Zhu et. al. zhu_insight_2023 showed that the effect of Hubbard interaction can be essential for the choice of magnetic phase in CrTe2. Also, the authors of Ref. gao_thickness_2021 emphasized that the on-site Coulomb repulsion changes the interlayer coupling to AFM even in the bulk; important role of the Coulomb repulsion for bilayer systems was also discussed in Ref. Interlayer . Previous theoretical works yao_control_2023 ; wu_-plane_2022 ; liu_structural_2022 ; li_tunable_2021 ; Interlayer used DFT$+U$ approach to account for correlation effects. However, for metallic systems this approach may be insufficient. To treat the effect of correlations we consider in the present study the DFT+DMFT approach. We use recently proposed method of evaluation of exchange interactions in paramagnetic phase MyJ to obtain unbiased estimate of exchange interactions, which account for correlation effects and consider also their temperature dependence. We compare the obtained results to those from DFT approach.
@@
The plan of the paper is the following. In Section II we describe used methods for magnetic properties of CrTe2 investigation. In Section III we present results for temperature and momentum dependence of magnetic susceptibilities, exchange interactions for monolayer, followed by study of bilayer and trilayer systems. In the end of the Section we consider bulk system to understand the limit of large number of layers. Finally, in Sect. IV we present Conclusions.
@@ II Methods and parameters
To obtain band structures we use DFT approach implemented in the Quantum Espresso QE package with ultrasoft pseudo-potentials from SSSP PBEsol Precision library PPP , supplemented by the maximally localized symmetry-adapted WanSym Wannier projection onto $3d$ states of Cr and $5p$ states of Te performed within Wannier90 package Wannier90 .
@@
For the bulk CrTe2 we choose the experimental lattice parameters freitas_ferromagnetism_2015 $a_{\rm bulk}=3.7887$, $c_{\rm bulk}=6.0955$. The momentum grid $18\times 18\times 18$ was chosen. For the single-layer CrTe2 we choose the lattice parameter of Ref. abuawwad_noncollinear_2022 $a=3.71$ and consider interlayer spacing $c=20$ to avoid overlap of the orbitals between the layers; the momentum grid $20\times 20\times 1$ was chosen. The vertical position of Te atoms was relaxed, and constituted the distance $1.55$ for the bulk and $1.62$ for the single-layer system from the Cr plane. For the bilayer (trilayer) CrTe2 we choose the lattice parameter of Ref. liu_structural_2022 $a=3.76$  ($a=a_{\rm bulk}$) and consider the unit cell of the vertical size $c=40$ (60 Å). The interlayer spacing was chosen equal to $c_{\rm bulk}$, Ref. sun_room_2020 . The relaxed vertical position of Te atoms constituted the distance $1.61$  ($1.56$) from the Cr plane for the outer (inner) atoms for the bilayer system and $1.58$  ($1.55$) for the trilayer system. The momentum grid was chosen $24\times 24\times 1$.
@@
The DMFT calculations of the self-energies, non-uniform susceptibilities and exchange interactions were performed within the continuous-time Quantum Monte Carlo (CT-QMC) method of the solution of impurity problem CT-QMC , realized in the iQIST software iQIST . In DMFT calculation we consider the basis, which diagonalises crystal field, and use the density-density interaction matrix, parameterized by Slater parameters $F^{0}$, $F^{2}$, and $F^{4}$, expressed through Hubbard $U$ and Hund $J_{H}$ interaction parameters according to ${F^{0}=U}$ and ${(F^{2}+F^{4})/14=J_{\rm H}}$, $F^{2}/F^{4}\simeq 0.63$ (see Ref. u_and_j, ). The Coulomb interaction parameters were chosen $U=2.8$ eV, $J_{H}=0.9$ eV. We use a double-counting correction in the around mean-field form AMF .
@@
To determine the exchange interactions we consider the effective Heisenberg model with the Hamiltonian $H=-(1/2)\sum_{{\bf q},rr^{\prime}}J^{rr^{\prime}}_{\bf q}{\mathbf{S}}^{r}_{\mathbf{q}}{\mathbf{S}}^{r^{\prime}}_{-{\mathbf{q}}}$, $\mathbf{S}^{r}_{\mathbf{q}}$ is the Fourier transform of static operators ${\mathbf{S}}_{ir}$ at the site $r$ of cell $i$, where the orbital-summed on-site static spin operators ${\mathbf{S}}_{ir}=\sum_{m}{\mathbf{S}}_{irm}$ The exchange interactions in DFT were estimated by magnetic force theorem (MFT) approach korotin_calculation_2015-1 ; liechtenstein_local_1987 ; git using electron Green’s functions, calculated in Wannier basis $G_{rr^{\prime},\sigma}^{mm^{\prime}}(\mathbf{k},i\nu_{n})=\left[(i\nu_{n}+\mu)I-H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}\right]_{rm,r^{\prime}m^{\prime}}^{-1}$, where $H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}$ is the Wannier Hamiltonian, $I$ is the identity matrix, and the inversion is performed in the site- and orbital space,
@@
The exchange interactions in paramagnetic phase in DMFT were calculated from the orbital-summed susceptibilities as MyJ ; MyCo ; Fe2C $\hat{J}_{\mathbf{q}}={\hat{\chi}}_{\mathrm{loc}}^{-1}-{\hat{\chi}}_{\mathbf{q}}^{-1}$, where hats stand for matrices $n\times n$ in the atom number in the unit cell, $\chi^{rr^{\prime}}_{\mathbf{q}}=-\langle\langle S_{\mathbf{q}}^{r,z}\mid S_{-\mathbf{q}}^{r^{\prime},z}\rangle\rangle_{\omega=0}$ is the matrix of the non-local static longitudinal susceptibilities, containing local vertex corrections, and obtained from the respective Bethe-Salpeter equation MyEDMFT ; OurRev with account of 60-80 fermionic Matsubara frequencies, including also corrections to finite frequency box MyEDMFT , $\chi^{rr^{\prime}}_{\mathrm{loc}}=-\left\langle\left\langle S_{i}^{r,z}\mid S_{i}^{r,z}\right\rangle\right\rangle_{\omega=0}\delta_{rr^{\prime}}$ is the diagonal matrix of local spin susceptibilities. The exchange interaction in real space is obtained by performing Fourier transform of $J_{\mathbf{q}}$.
@@ III Results
Figure 1: (Color online) Band structure (a) and DFT density of states (b) of the single-layer CrTe2 in paramagnetic phase. The inset in (a) shows the DFT Fermi surface; the arrow shows the nesting vector which is close to ${\mathbf{Q}}_{K}=\overrightarrow{\Gamma K}$.
@@
The DFT band structure of single-layer CrTe2 in the paramagnetic state is shown in Fig. 1. In an agreement with the earlier study sun_room_2020 , the density of states contains weak peak at the Fermi level, which originates from the flat parts of the dispersion. The corresponding Fermi surface contains only small pockets near $K$, $K^{\prime}$, and $M$ points of the Brillouin zone.
@@
The filling of d-orbitals of chromium, obtained in DFT+DMFT calculations for the single layer compound CrTe2, is $n_{\rm Cr}=4.83$, and, therefore, sufficiently close to half filling, which provides rather strong correlation effects. In Fig. 2 we show imaginary parts of the electronic self-energy for combinations of orbitals, which diagonalize crystal field (the rotation angle $\theta\simeq\pi/3$). One can see that the for some of the states (those which are close to the Fermi level) the damping of electronic excitations is particularly large, and the self-energy for some of the combinations has a non-quasiparticle form with $\partial{\rm Im}\Sigma/\partial\nu>0$.
@@
Figure 2: Imaginary part of the self-energy of various orbital combinations of single-layer (blue solid symbols) and bilayer (green open symbols) CrTe2 at the imaginary frequency axis in DFT+DMFT approach at $\beta=18$ eV-1.
@@
Figure 3: Momentum dependence of the orbital-summed susceptibility of the single-layer CrTe2 in DFT+DMFT approach at $\beta=7$ eV-1.
@@
The obtained magnetic susceptibility at $\beta=7$ eV-1 is rather large (see Fig. 3) and has a maximum at the wave vector $\mathbf{Q}_{K}=(4\pi,0)/(3a)$, corresponding to the tendency to the $120\degree$ incommensurate magnetic order. Despite strong incoherence of electronic excitations, this tendency is likely related to the nesting of the pockets of the Brillouin zone, as shown in the inset of Fig. 1a. The susceptibility does not show even local maximum near the $\Gamma$ point, which implies an absence of pronounced tendency towards ferromagnetic order.
@@
Figure 4: (Color online) Inverse uniform and staggered, ${\mathbf{Q}}={\mathbf{Q}}_{K}$ (a) and local (b) spin susceptibility of the single-layer CrTe2 in DFT+DMFT approach. Dashed lines show extrapolation to the low-temperature region.
@@
The temperature dependence of the inverse uniform and staggered magnetic susceptibilities is shown in Fig. 4. The corresponding DMFT Neel temperature, obtained by the extrapolation of the inverse staggered susceptibility, $T^{\rm DMFT,1{\text{-}}layer}_{N}\simeq 300K$. At the same time, the inverse uniform susceptibility is extrapolated to the small negative Weiss temperature, showing absence of the ferromagnetic order, in agreement with the conclusion from the momentum dependence of the non-uniform susceptibility. The inverse local susceptibility (see Fig. 4b) is almost linear in temperature, showing well formed local magnetic moments with negligibly small Kondo temperature, estimated as an offset of linear extrapolation from zero. The local magnetic moment, obtained from the slope of the local (staggered) susceptibility $\mu_{\rm loc}^{2}=32.4\mu_{B}^{2}$ ($\mu^{2}=31.7\mu_{B}^{2}$), corresponding to the local effective spin $S_{\rm eff}\simeq 2.4$ defined by $\mu^{2}=(g\mu_{B})^{2}S_{\rm eff}(S_{\rm eff}+1)$, $g=2$ is the spin $g$-factor.
@@
Figure 5: (Color online) Exchange interaction in single-layer CrTe2 at $\beta=20$ eV-1. Red solid line (squares) corresponds to DFT+DMFT approach in paramagnetic phase, blue dashed line (filled triangles) is the result of the DFT approach in FM phase. For comparison, the result of DFT approach at $\beta=10$ eV-1 is shown by dot-dashed violet line with open triangles
@@
Magnetic exchange interactions of single-layer CrTe2 at $\beta=20$ eV-1 are shown in Fig. 5. The maximum of the exchange interaction in both DFT and DFT+DMFT approaches is also found at the wave vector ${\bf Q}_{K}$, which corresponds to the 120°$\degree$ antiferromagnetic order. Locally FM order is stable in DFT due to local maximum at the $\Gamma$ point with $J_{0}>0$ and unstable in DMFT, but in both cases $\Gamma$ point is not a global maximum, which corresponds to an instability of ferromagnetism in Heisenberg model. We find the peak of exchange interaction $J_{\mathbf{q}}$ at $\mathbf{q}=\mathbf{Q}_{K}$ in DFT approach sharply increases with decrease of temperature (see the results of $\beta=10$ eV-1 in Fig. 5 for comparison), due to nesting of the Fermi surface pockets with the wave vector close to $\mathbf{Q}_{K}$ (see Appendix A). At the same time, the exchange interactions in DFT+DMFT approach are quite weakly temperature dependent, because of the damping of electronic excitations. We also note that in the ferromagnetic ordered phase of DFT we obtain magnetic moment of chromium sites $\mathfrak{m}=2.86\mu_{B}$, which corresponds to effective spin $S_{\rm ord}=\mathfrak{m}/(g\mu_{B})=1.48$. This reflects different spin state of chromium in the ferro- and paramagnetic phase: while the chromium occupation is close in two approaches ($n_{\rm Cr}=4.67$ in DFT), the effective spin difference $\Delta S=S_{\rm eff}-S_{\rm ord}\simeq 1$ corresponds to only partial magnetization of orbital states. In particular, we find strongest magnetization contribution $0.8\mu_{B}$/f.u. from $d_{3z^{2}-r^{2}}$ state and weakest contribution $0.4\mu_{B}$/f.u. from each of the $d_{xz,yz}$ states.
@@
DFT Energies, this work $-2.03$ $1.53$ $-0.43$
@@
Table 1: Exchange interactions (in meV) in monolayer CrTe2 for nearest ($J_{1}$), next nearest ($J_{2}$) and next to next nearest ($J_{3}$) neighbours. DFT MFT and DMFT exchange interactions are calculated at $\beta=20$ eV-1.

### v3

@@ Abstract
We investigate magnetic properties of CrTe2 within the density functional theory (DFT) approach in ferromagnetic phase and the combination of the DFT and the dynamical mean field theory (DFT+DMFT) approach in paramagnetic phase. We show that a few layer CrTe2 possesses well formed local magnetic moments. In the monolayer CrTe2 we find the most preferable antiferromagnetic exchange with the 120°$\degree$ antiferromagnetic structure. In the bilayer and trilayer systems, electronic correlations in DFT+DMFT approach yield ferromagnetic exchange interaction within each layer, but the interaction between the layers is antiferromagnetic, such that alternation of the direction of magnetization of the layers is expected. In bulk CrTe2 we find the tendency to ferromagnetic order at low temperature, but with increase of temperature antiferromagnetic correlations between the layers dominate. Determination of the critical number of layers at which the interlayer antiferromagnetic order changes to the ferromagnetic likely requires consideration of the non-local Coulomb interactions.
@@ I Introduction
Atomically thin two-dimensional (2D) van der Waals (vdW) materials have attracted significant attention over the past decade due to their exceptional properties Rev1 ; Rev2 ; Rev3 ; Rev4 ; Rev5 ; Rev6 . In particular, they can posses large magnetic moments, substantial Curie temperature, comparable to room temperature, and strong anisotropy. With these properties, 2D magnetic materials are promising for using in the fields of spintronics Rev2 and memory storage devices Rev5 . Currently, many 2D ferromagnetic (FM) materials have been reported; however, a key challenge remains: achieving ferromagnetism at room temperature. According to the Mermin-Wagner theorem, a long-range magnetic order cannot exist in two dimensions in isotropic systems at finite temperatures due to thermal fluctuations. However, it is possible to circumvent this limitation by magnetic anisotropy.
@@
Recently, chromium ditelluride CrTe2 has been suggested to exhibit ferromagnetic order even up to room temperature, making it an attractive candidate for further research in this field. Bulk CrTe2 was shown freitas_ferromagnetism_2015 to posses FM order with $T_{C}=330$ K and in-plane magnetic anisotropy. Recent years several groups obtained thin CrTe2 flakes via mechanical exfoliation purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , chemical vapor deposition meng_anomalous_2021 and molecular beam epitaxy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; liu_wafer-scale_2023 ; zheng_two-dimensional_2023 , showing stable FM order with $T_{C}\sim 200$ K. Although thick films were shown to have an in-plane magnetic anisotropy purbawati_-plane_2020 ; sun_room_2020 ; fabre_characterization_2021 , the films thinner 10 monolayers show out-of-plane magnetic anisotropy sun_ferromagnetism_2021 ; zhang_room-temperature_2021 ; meng_anomalous_2021 . The authors of Refs. zhang_room-temperature_2021 ; meng_anomalous_2021 ; xian_spin_2022 ; liu_wafer-scale_2023 ; wang_strain-_2024 succeeded fabricating monolayer devices. At the same time, the results of the magnetic state of monolayers remain controversial. Experimentally both, FM zhang_room-temperature_2021 ; meng_anomalous_2021 ; liu_wafer-scale_2023 and zigzag antiferromagnetic (AFM-zz) xian_spin_2022 states were reported.
@@
Theoretically, the density functional theory (DFT) approach was used to study various magnetic orders and exchange interactions in CrTe2. In particular, FM lv_strain-controlled_2015 ; li_tunable_2021 ; liu_structural_2022 ; yang_tunable_2021 ; yao_control_2023 , AFM gao_thickness_2021 ; AFM2 ; abuawwad_noncollinear_2022 , charge density wave otero_fumega_controlled_2020 and incommensurate abuawwad_noncollinear_2022 states were considered as possible candidates in monolayer CrTe2. It was suggested in Refs. wu_-plane_2022 ; zhu_insight_2023 ; gao_thickness_2021 ; lv_strain-controlled_2015 ; AFM2 that relatively small strain can cause the transition from AFM-zz to FM, which may explain different kinds of magnetic order observed in experiment. Furthermore, it was shown in Ref. gao_thickness_2021 that the intralayer coupling of 2 to 4 layer CrTe2 is still antiferromagnetic in DFT, similar to that of monolayer CrTe2, and changes to ferromagnetic for a larger number of layers. At the same time, the interlayer exchange interaction was also found to change from AFM to FM between 4 and 5 layers. The change of the magnetic orders was related by the authors of Ref. gao_thickness_2021 to the change of the lattice constants from monolayer to bulk. However, later study of Ref. wu_-plane_2022 has shown that at least in the bilayer system the in-plane antiferromagnetic order is stable in DFT for a wide range of lattice constants.
@@
At the same time, since Cr is an open almost half-filled $d$ shell metal, it possesses strong correlations, which may affect preferable magnetic state. Zhu et. al. zhu_insight_2023 showed that the effect of Hubbard interaction can be essential for the choice of magnetic phase in CrTe2. Also, the authors of Ref. gao_thickness_2021 emphasized that the on-site Coulomb repulsion changes the interlayer coupling to AFM even in the bulk; important role of the Coulomb repulsion for bilayer systems was also discussed in Ref. Interlayer . Previous theoretical works yao_control_2023 ; wu_-plane_2022 ; liu_structural_2022 ; li_tunable_2021 ; Interlayer used DFT$+U$ approach to account for correlation effects. However, for metallic systems this approach may be insufficient. To treat the effect of correlations, we consider in the present study the DFT+DMFT approach. We use recently proposed method of evaluation of exchange interactions in the paramagnetic phase MyJ to obtain an unbiased estimate of exchange interactions, which account for correlation effects and also consider their temperature dependence. We compare the results obtained with those from the DFT approach.
@@
The plan of the paper is the following. In Section II we describe methods used for the magnetic properties of CrTe2 investigation. In Section III we present results for temperature and momentum dependence of magnetic susceptibilities, exchange interactions for monolayer systems, followed by a study of bilayer and trilayer systems. In the end of the Section we consider bulk system to understand the limit of large number of layers. Finally, in Sect. IV we present Conclusions.
@@ II Methods and parameters
To obtain band structures, we use the DFT approach implemented in the Quantum Espresso QE package with ultrasoft pseudopotentials from the SSSP PBEsol Precision library PPP , supplemented by the maximally localized symmetry-adapted WanSym Wannier projection on $3d$ Cr states and $5p$ Te states performed within the Wannier90 package Wannier90 .
@@
For bulk CrTe2 we choose the experimental lattice parameters freitas_ferromagnetism_2015 $a_{\rm bulk}=3.7887$, $c_{\rm bulk}=6.0955$. The momentum grid $18\times 18\times 18$ was chosen. For the monolayer CrTe2 we choose the lattice parameter of Ref. abuawwad_noncollinear_2022 $a=3.71$ and consider interlayer spacing $c=20$ to avoid overlap of the orbitals between the layers; the momentum grid $20\times 20\times 1$ was chosen. The vertical position of Te atoms was relaxed and constituted the distance $1.55$ for the bulk and $1.62$ for the monolayer system from the Cr plane. For the bilayer (trilayer) CrTe2 we choose the lattice parameter of Ref. liu_structural_2022 $a=3.76$  ($a=a_{\rm bulk}$) and consider the unit cell of vertical size $c=40$ (60 Å). The interlayer spacing was chosen equal to $c_{\rm bulk}$, Ref. sun_room_2020 . The relaxed vertical position of Te atoms constituted the distance $1.61$  ($1.56$) from the Cr plane for the outer (inner) atoms for the bilayer system and $1.58$  ($1.55$) for the trilayer system. The momentum grid was chosen $24\times 24\times 1$.
@@
DMFT calculations of self-energies, non-uniform susceptibilities, and exchange interactions were performed within the continuous-time Quantum Monte Carlo (CT-QMC) method of the solution of impurity problem CT-QMC , realized in the iQIST software iQIST . In DMFT calculation we consider the basis, which diagonalises crystal field, and use the density-density interaction matrix, parameterized by Slater parameters $F^{0}$, $F^{2}$, and $F^{4}$, expressed through Hubbard $U$ and Hund $J_{H}$ interaction parameters according to ${F^{0}=U}$ and ${(F^{2}+F^{4})/14=J_{\rm H}}$, $F^{2}/F^{4}\simeq 0.63$ (see Ref. u_and_j, ). The Coulomb interaction parameters were chosen $U=2.8$ eV, $J_{H}=0.9$ eV. We use a double-counting correction in the around mean-field form AMF .
@@
To determine exchange interactions, we consider the effective Heisenberg model with the Hamiltonian $H=-(1/2)\sum_{{\bf q},rr^{\prime}}J^{rr^{\prime}}_{\bf q}{\mathbf{S}}^{r}_{\mathbf{q}}{\mathbf{S}}^{r^{\prime}}_{-{\mathbf{q}}}$, $\mathbf{S}^{r}_{\mathbf{q}}$ is the Fourier transform of static operators ${\mathbf{S}}_{ir}$ at the site $r$ of cell $i$, where the orbital-summed static spin operators ${\mathbf{S}}_{ir}=\sum_{m}{\mathbf{S}}_{irm}$ The exchange interactions in DFT were estimated by magnetic force theorem (MFT) approach korotin_calculation_2015-1 ; liechtenstein_local_1987 ; git using electron Green’s functions, calculated in Wannier basis $G_{rr^{\prime},\sigma}^{mm^{\prime}}(\mathbf{k},i\nu_{n})=\left[(i\nu_{n}+\mu)I-H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}\right]_{rm,r^{\prime}m^{\prime}}^{-1}$, where $H^{rm,r^{\prime}m^{\prime}}_{\mathbf{k},\sigma}$ is the Wannier Hamiltonian, $I$ is the identity matrix, and the inversion is performed in the site- and orbital space,
@@
The exchange interactions in paramagnetic phase in DMFT were calculated from the orbital-summed susceptibilities as MyJ ; MyCo ; Fe2C ; CrO2 $\hat{J}_{\mathbf{q}}={\hat{\chi}}_{\mathrm{loc}}^{-1}-{\hat{\chi}}_{\mathbf{q}}^{-1}$, where hats stand for matrices $n\times n$ in the atom number in the unit cell, $\chi^{rr^{\prime}}_{\mathbf{q}}=-\langle\langle S_{\mathbf{q}}^{r,z}\mid S_{-\mathbf{q}}^{r^{\prime},z}\rangle\rangle_{\omega=0}$ is the matrix of the non-local static longitudinal susceptibilities, containing local vertex corrections, and obtained from the respective Bethe-Salpeter equation MyEDMFT ; OurRev with account of 60-80 fermionic Matsubara frequencies, including also corrections to finite frequency box MyEDMFT (at low temperatures we also use the method of Ref. Loon ), $\chi^{rr^{\prime}}_{\mathrm{loc}}=-\left\langle\left\langle S_{i}^{r,z}\mid S_{i}^{r,z}\right\rangle\right\rangle_{\omega=0}\delta_{rr^{\prime}}$ is the diagonal matrix of local spin susceptibilities. The exchange interaction in real space is obtained by performing the Fourier transform of $J_{\mathbf{q}}$.
@@ III Results
Figure 1: (Color online) Band structure (a) and DFT density of states (b) of the monolayer CrTe2 in paramagnetic phase. The inset in (a) shows the DFT Fermi surface; the arrow shows the nesting vector which is close to ${\mathbf{Q}}_{K}=\overrightarrow{\Gamma K}$.
@@
The DFT band structure of the monolayer CrTe2 in the paramagnetic state is shown in Fig. 1. In agreement with the earlier study sun_room_2020 , the density of states contains a weak peak at the Fermi level, which originates from the flat parts of the dispersion. The corresponding Fermi surface contains only small pockets near $K$, $K^{\prime}$, and $M$ points of the Brillouin zone.
@@
The filling of the chromium d orbitals, obtained in DFT+DMFT calculations for the monolayer compound CrTe2, is $n_{\rm Cr}=4.83$, and therefore sufficiently close to half filling, which provides rather strong correlation effects. In Fig. 2 we show imaginary parts of the electronic self-energy for combinations of orbitals, which diagonalize crystal field (the rotation angle $\theta\simeq\pi/3$). One can see that the for some of the states (those which are close to the Fermi level) the damping of electronic excitations is particularly large, and the self-energy for some of the combinations has a non-quasiparticle form with $\partial{\rm Im}\Sigma/\partial\nu>0$.
@@
Figure 2: Imaginary part of the self-energy of various orbital combinations of monolayer (blue solid symbols) and bilayer (green open symbols) CrTe2 at the imaginary frequency axis in DFT+DMFT approach at $\beta=18$ eV-1 ($T=645$ K).
@@
Figure 3: Momentum dependence of the orbital-summed susceptibility of the monolayer CrTe2 in DFT+DMFT approach at $\beta=7$ eV-1 ($T=1660$ K).
@@
The magnetic susceptibility obtained at $\beta=7$ eV-1 ($T=1660$ K) is rather large (see Fig. 3) and has a maximum at the wave vector $\mathbf{Q}_{K}=(4\pi,0)/(3a)$, corresponding to the tendency to the $120\degree$ incommensurate magnetic order. Despite the strong incoherence of electronic excitations, this tendency is likely related to the nesting of the pockets of the Brillouin zone, as shown in the inset of Fig. 1a. The susceptibility does not show even a local maximum near the $\Gamma$ point, which implies an absence of pronounced tendency towards ferromagnetic order.
@@
Figure 4: (Color online) Inverse uniform and staggered, ${\mathbf{Q}}={\mathbf{Q}}_{K}$ (a) and local (b) spin susceptibility of the monolayer CrTe2 in DFT+DMFT approach. Dashed lines show extrapolation to the low-temperature region.
@@
The temperature dependence of the inverse uniform and staggered magnetic susceptibilities is shown in Fig. 4. The corresponding DMFT Neel temperature, obtained by extrapolating the inverse staggered susceptibility, $T^{\rm DMFT,1{\text{-}}layer}_{N}\simeq 300K$. At the same time, the inverse uniform susceptibility is extrapolated to the small negative Weiss temperature, showing absence of the ferromagnetic order, in agreement with the conclusion from the momentum dependence of the nonuniform susceptibility. The inverse local susceptibility (see Fig. 4b) is almost linear in temperature, showing well-formed local magnetic moments with negligibly small Kondo temperature, estimated as an offset of linear extrapolation from zero. The local magnetic moment, obtained from the slope of the local (staggered) susceptibility $\mu_{\rm loc}^{2}=32.4\mu_{B}^{2}$ ($\mu^{2}=31.7\mu_{B}^{2}$), corresponding to the local effective spin $S_{\rm eff}\simeq 2.4$ defined by $\mu^{2}=(g\mu_{B})^{2}S_{\rm eff}(S_{\rm eff}+1)$, $g=2$ is the spin $g$-factor.
@@
Figure 5: (Color online) Exchange interaction in monolayer CrTe2 at $\beta=20$ eV-1 ($T=580$ K). Red solid line (squares) corresponds to DFT+DMFT approach in paramagnetic phase, blue dashed line (filled triangles) is the result of the DFT approach in FM phase. For comparison, the result of DFT approach at $\beta=10$ eV-1 ($T=1160$ K) is shown by dot-dashed violet line with open triangles
@@
Magnetic exchange interactions of the monolayer CrTe2 at $\beta=20$ eV-1 ($T=580$ K) are shown in Fig. 5. The maximum of exchange interaction in both, the DFT and DFT + DMFT approaches is also found at the wave vector ${\bf Q}_{K}$, which corresponds to the 120°$\degree$ antiferromagnetic order. FM order is locally stable in DFT due to the local maximum at the $\Gamma$ point with $J_{0}>0$ and unstable in DMFT, but in both cases the $\Gamma$ point is not a global maximum, which corresponds to an instability of ferromagnetism in Heisenberg model. We find that the peak of exchange interaction $J_{\mathbf{q}}$ at $\mathbf{q}=\mathbf{Q}_{K}$ in DFT approach sharply increases with a decrease of temperature (see the results of $\beta=10$ eV-1 in Fig. 5 for comparison), due to nesting of the Fermi surface pockets with the wave vector close to $\mathbf{Q}_{K}$ (see Appendix A). At the same time, the exchange interactions in the DFT+DMFT approach are quite weakly temperature-dependent because of the damping of electronic excitations. We also note that in the ferromagnetic ordered phase of DFT we obtain the magnetic moment of chromium sites $\mathfrak{m}=2.86\mu_{B}$, which corresponds to the effective spin $S_{\rm ord}=\mathfrak{m}/(g\mu_{B})=1.48$. This reflects different spin states of chromium in the ferro- and paramagnetic phases: while the chromium occupation is close in two approaches ($n_{\rm Cr}=4.67$ in DFT), the effective spin difference $\Delta S=S_{\rm eff}-S_{\rm ord}\simeq 1$ corresponds to only partial magnetization of orbital states. In particular, we find the strongest magnetization contribution $0.8\mu_{B}$ / f.u. from the $d_{3z^{2}-r^{2}}$ state and the weakest contribution $0.4\mu_{B}$ / f.u. from each of the $d_{xz,yz}$ states.
@@
DFT Energies, this work ${-1.52}$ 1.021.02{1.02}1.02 ${-0.17}$
@@
Table 1: Exchange interactions (in meV) in monolayer CrTe2 for nearest ($J_{1}$), next nearest ($J_{2}$) and next to next nearest ($J_{3}$) neighbours. DFT MFT and DMFT exchange interactions are calculated at $\beta=20$ eV-1 ($T=580$ K).